# Edge Agentic AI Framework for Autonomous Network Optimisation in O-RAN

### v2

@@ Abstract
ORAN, Agentic AI, Edge Agent, GenAI, 6G network, Autonomous Network Optimisation.
@@ I Introduction
The Open Radio Access Network (O-RAN) architecture enables intelligent network control through RAN Intelligent Controllers (RICs) that support components such as xApps and rApps, which are modular applications that can dynamically manage radio resources. O-RAN employs a disaggregated architecture that divides traditional gNodeB functionality across three open units: the Radio Unit (RU), Distributed Unit (DU), and Centralised Unit (CU). This allows mobile network operators (MNOs) to optimise performance, leverage cloud elasticity, and meet the diverse demands of different applications and services within the O-RAN framework [2]. Individual components can be invoked manually by the MNO or, more implicitly, through Agentic AI frameworks to meet performance goals [3, 4]. However, integrating LLM-based Agentic AI agents in O-RAN environments presents unique challenges due to task specific operational timescales and control loop requirements. These include ensuring trustworthy decision-making under strict latency and safety constraints, dealing with limited compute and memory resources at the edge, and managing the complexity of large-scale agent coordination.
To address these challenges, we present an Agentic AI framework within edge-based RIC for network optimisation, as shown in Fig. 1. We present a novel Agentic AI framework that incorporates two key innovations: (i) a persona-based agent model and (ii) embedded search tools along with Long Short-Term Memory (LSTM) xApp for predictive traffic analytics. The persona-based architecture allows agents to simulate specialised roles, such as outage mitigator or energy optimiser, drawing from recent advances in contextual agent design for reasoning, dialogue, and control [5, 6, 7]. Meanwhile, the integrated LSTM model continuously forecasts network load and signal quality trends over future time horizons. By embedding the LSTM directly into the RIC environment, the framework becomes capable of proactive outage prevention, fine-grained power control, and early anomaly detection—essential capabilities for near real-time operations at the edge.
@@ II Related work
Contemporary wireless networks leverage diverse AI technologies for network optimisation and management. LLMs, including GPT-4, LLaMA, and domain-specific BERT variants, are being deployed by industry leaders such as NVIDIA, Ericsson, SoftBank, AT&\&&T, Nokia, and Vodafone for automated network operations [8, 9, 10]. Autonomous AI agents capable of independent decision-making are being operationalised for complex network orchestration tasks [11, 12]. Within the RAN Intelligent Controller (RIC) framework, reinforcement learning agents and hierarchical planners orchestrate xApps and rApps, forming the backbone of intelligent RAN operations [13, 14]. Advanced agentic frameworks supporting multi-step planning and complex reasoning are under active investigation in both academic and industrial research contexts [5, 15]. LLMs are trained or fine-tuned on telecom-relevant datasets, including standards documentation and troubleshooting logs [16]. The work by Dev et al. [17] presents a forward-looking framework for integrating Agentic AI into 5G architectures, offering benefits such as simplified network operations through user/control plane separation, dynamic service orchestration via serverless computing, and enhanced adaptability using autonomous cognitive agents and neural radio protocol stacks. It highlights substantial gains in energy efficiency, latency reduction, and robustness, particularly in V2X scenarios.
@@
In prediction and classification, LLMs contribute to tasks such as traffic forecasting, signal processing, and security threat detection [24]. Pre-trained models like BERT and GPT have been adapted to classify encrypted traffic and identify network anomalies [25].
@@ III-C AI Agent Decision Framework
  • Tier 1 Critical (SINR $<\gamma_{1}$ dB): Apply emergency power boost (increase transmission power by up to 3 dB) and allocate additional Physical Resource Blocks (PRBs) to maintain link stability and prevent imminent outage.
  • Tier 2 Risk ($\gamma_{1}\leq$ SINR $<\gamma_{2}$ dB): Initiate preventive adjustments (+1 to +3 dB) to avoid potential degradation in link quality.
  • Tier 3 Optimization ($\gamma_{2}\leq$ SINR $<\gamma_{3}$ dB): Perform fine-grained tuning ($\pm$1 to $\pm$2 dB) to balance performance with energy consumption.
@@
where $\gamma_{1}$ denotes the critical network threshold, set to 15 dB in this work to meet Ultra-Reliable Low Latency (URLLC) network needs. The remaining thresholds are defined as follows: $\gamma_{2}$ = 18 dB, $\gamma_{3}$ = 20 dB, and $\gamma_{4}$ = 25 dB.
@@ III-E Key Performance Indicators
Outage Rate: Percentage of time SINR falls below the 15 dB threshold for Ultra-Reliable Low Latency (URLLC):
@@ IV-B Edge Deployment Advantages and Real-time Performance
Critical performance evaluation during high-stress event scenarios demonstrated 100% outage prevention success rate compared to 91.6% (traditional) and 96.7% (standard LLM-based baseline). The model maintained a coefficient of variation below 0.15 across all primary metrics during the 15 validation event days, demonstrating operational consistency.
While our results demonstrate significant improvements with strong statistical support, some limitations warrant consideration for future development. Scalability across diverse network topologies and operator environments needs further validation through large-scale deployments beyond the current minute-resolution 2-month simulation time frame.

### v3

@@ Abstract
O-RAN, Agentic AI, Edge Agent, GenAI, 6G Networks, Anomaly Detection, Network Optimisation.
@@ I Introduction
The Open Radio Access Network (O-RAN) architecture enables intelligent network control through RAN Intelligent Controllers (RICs) that support components such as xApps and rApps, which are modular applications that can dynamically manage radio resources. O-RAN employs a disaggregated architecture that divides traditional gNodeB functionality across three open units: the Radio Unit (RU), Distributed Unit (DU), and Centralised Unit (CU). This allows mobile network operators (MNOs) to optimise performance, leverage cloud elasticity, and meet the diverse demands of different applications and services within the O-RAN framework [2]. Individual components can be invoked manually by the MNO or, more implicitly, through Agentic AI frameworks to meet performance goals [3, 4]. However, integrating LLM-based models in O-RAN environments presents unique challenges due to task specific operational timescales and control loop requirements. These include ensuring trustworthy decision-making under strict latency and safety constraints, dealing with limited compute and memory resources at the edge, and managing the complexity of large-scale agent coordination.
@@
To address these challenges, we present an Agentic AI framework within edge-based RIC for network optimisation, as shown in Fig. 1. We present a novel Agentic AI framework that incorporates two key innovations: (i) a persona-based agent model and (ii) embedded search tools along with Long Short-Term Memory (LSTM) xApp for predictive traffic analytics. The persona-based architecture allows agents to simulate specialised roles, such as outage mitigator or energy optimiser, drawing from recent advances in contextual agent design for reasoning, dialogue, and control [5, 6, 7]. Meanwhile, the integrated LSTM model continuously forecasts network load and signal quality trends over future time horizons. LSTM integration within the RIC enables proactive outage prevention, fine-grained power control, and early anomaly detection for near real-time edge operations.
@@ II Related work
Contemporary wireless networks leverage diverse AI technologies for network optimisation and management. LLMs, including GPT-4, LLaMA, and domain-specific BERT variants, are being deployed by industry leaders such as NVIDIA, Ericsson, SoftBank, AT&\&T, Nokia, and Vodafone for automated network operations [8, 9, 10]. Autonomous AI agents capable of independent decision-making are being operationalised for complex network orchestration tasks [11, 12]. Within the RAN Intelligent Controller (RIC) framework, reinforcement learning agents and hierarchical planners orchestrate xApps and rApps, forming the backbone of intelligent RAN operations [13, 14]. Advanced agentic frameworks supporting multi-step planning and complex reasoning are under active investigation in both academic and industrial research contexts [5, 15]. LLMs are trained or fine-tuned on telecom-relevant datasets, including standards documentation and troubleshooting logs [16]. The work by Dev et al. [17] presents a forward-looking framework for integrating Agentic AI into 5G architectures, offering benefits such as simplified network operations through user/control plane separation, dynamic service orchestration via serverless computing, and enhanced adaptability using autonomous cognitive agents and neural radio protocol stacks. It highlights substantial gains in energy efficiency, latency reduction, and robustness, particularly in V2X scenarios.
@@
In prediction and classification, LLMs contribute to tasks such as traffic forecasting, signal processing, and security threat detection. Pre-trained models like BERT and GPT have been adapted to classify encrypted traffic and identify network anomalies [25].
@@ III-C AI Agent Decision Framework
  • Tier 1 Critical (SINR $\gamma_{1}$ dB): Apply emergency power boost (increase transmission power by up to 3 dB) and allocate additional Physical Resource Blocks (PRBs) to maintain link stability and prevent imminent outage.
  • Tier 2 Risk ($\gamma_{1}\leq$ SINR $\gamma_{2}$ dB): Initiate preventive adjustments (+1 to +3 dB) to avoid potential degradation in link quality.
  • Tier 3 Optimization ($\gamma_{2}\leq$ SINR $\gamma_{3}$ dB): Perform fine-grained tuning ($\pm$1 to $\pm$2 dB) to balance performance with energy consumption.
@@
where $\gamma_{1}$ denotes the critical network threshold, set to 15 dB in this work to satisfy high QoS requirements. The remaining thresholds are defined as follows: $\gamma_{2}$ = 18 dB, $\gamma_{3}$ = 20 dB, and $\gamma_{4}$ = 25 dB.
@@ III-D LSTM xApp Integration into the Agentic AI Framework
The LSTM model employs a two-layer LSTM architecture (128, 64 units) followed by dense layers (32, 16 units). Training parameters include 500 epochs, batch size 32, learning rate 0.001, dropout 0.2, early stopping 10, and Adam optimiser.
@@ III-E Key Performance Indicators
Outage Rate: Percentage of time SINR falls below the threshold (i.e., $\gamma_{1}$ dB):
@@ IV-B Edge Deployment Advantages and Real-time Performance
Critical performance evaluation during high-stress event scenarios (i.e., high traffic demand) demonstrates 100% outage prevention success rate in compared to 91.6% (traditional) and 96.7% (standard LLM-based baseline). The model maintained a coefficient of variation below 0.15 across all metrics during the 15 validation event days, showing operational consistency.
@@
While our results demonstrate significant improvements with strong statistical support, some limitations warrant consideration for future development. Scalability across diverse network topologies and operator environments needs further validation through large-scale deployments.
@@ V Conclusion
Acknowledgment
This research was funded by EPSRC CHEDDAR (EP/X040518/1), UKRI Grant EP/X039161/1, and MSCA Horizon EU Grant 101086218. Additionally, this research was supported by the UKRI Funding Service under Award UKRI851: Strategic Decision-Making and Cooperation among AI Agents: Exploring Safety and Governance in Telecom.